# An Entropy-based Text Watermarking Detection Method

### v3

@@ Abstract
Currently, text watermarking algorithms for large language models (LLMs) can embed hidden features to texts generated by LLMs to facilitate subsequent detection, thus alleviating the problem of misuse of LLMs. Although the current text watermarking algorithms perform well in most high-entropy scenarios, its performance in low-entropy scenarios still needs to be improved. In this work, we proposed that the influence of token entropy should be fully considered in the watermark detection process, that is, the weight of each token during watermark detection should be customized according to its entropy, rather than setting the weights of all tokens to the same value as in previous methods. Specifically, we proposed an Entropy-based Text Watermark Detection (EWD) that gives higher-entropy tokens higher influence weights during watermark detection, so as to better reflect the degree of watermarking. Furthermore, the proposed detection process is training-free and fully automated. In the experiment, we found that our method can achieve better detection performance in low-entropy scenarios, and our method is also general and can be applied to texts with different entropy distributions. Our code and data is available on https://github.com/luyijian3/EWD. Additionally, our algorithm could be accessed through MarkLLMPan et al. (2024).
@@
Yijian LU1, Aiwei Liu2 , Dianzhi Yu1 , Jingjing Li1 , Irwin King1 1The Chinese University of Hong Kong 2Tsinghua University luyijian@link.cuhk.edu.hk, liuaw20@mails.tsinghua.edu.cn king@cse.cuhk.edu.hk
@@ 1 Introduction
Figure 1: Watermarking algorithm has limited effect on a low-entropy token as compared to a high-entropy token. Therefore, it is almost random for a low-entropy token to be green. In a code generation scenario, most of the tokens are low-entropy, resulting in less green tokens and a small detection z-score.
@@
The advancements in LLMs have enabled it to generate high-quality outputs indistinguishable from humans, and also achieve better performance on various real-world generation tasks. However, this gives rise to the potential risk of misuse. Specifically, LLM-generated assignments, including essays, code, and reports, pose a threat to academic integrity. Misusing LLMs to generate fake news can also lead to negative societal consequences Megías et al. (2021). Therefore, algorithms capable of effectively marking and identifying machine-generated contents have become essential.
The text watermarking algorithm for LLMs can embed hidden features that can be detected later during the text generation process of LLMs, thus effectively alleviating the problem of misuse of LLMs. The current LLM watermarking technology is usually to embed watermark information in the process of generating tokens during LLM inference. For example, KGW Kirchenbauer et al. (2023a) divides the model’s vocabulary into two lists, green and red, and increases the logit value of tokens in the green list. By modifying the token distributions, it imposes a preference on the green list to the model and leads the model to sample most tokens from the green list. In the detection process, the degree of watermark is usually judged by the number of green tokens in the text.
Although current text watermarking algorithms perform well in most high-entropy scenarios, its performance in low-entropy scenarios such as code generation or formatted document generation still needs to be improved. The reason behind this is that low-entropy tokens usually have stronger determinism, and the modified logits can hardly affect the probability of sampling the tokens from the green list. At the same time, if logits is forcefully modified, the quality of the watermarked text is easier to degrade Tu et al. (2023).
In figure 1, we illustrate the green-list ratio in tokens of three different entropy levels. A token with higher entropy will have a significantly larger probability to be sampled from a green list. On the contrary, it is almost random to see a green token in low-entropy tokens if the green list size is half of the model vocabulary size.
@@
When the generated text contains a large number of low-entropy tokens, tokens that can hardly be affected by the modified logits, the detection results of the watermark detector are more easily affected by the states (colors) of these low-entropy tokens even when the watermark generator managed to sample all high-entropy tokens from the green list. For example, the vast low-entropy tokens in a machine-generated text could all appear to be red, and cause the detector to give a low detection score, thus leading to more misjudgments. To address this problem, Lee et al. (2023) proposed a selective watermark detection method named SWEET. Specifically, they only detect tokens that exceed a certain entropy threshold. However, this method has some problems. First, their method requires a manually-set entropy threshold by analyzing samples of code written by humans. At the same time, their method also does not consider the distribution of token entropy, that is, their method treats all tokens that exceed the threshold as the same, while in fact the entropy of these tokens may vastly varies.
@@
To address the aforementioned problems on low-entropy watermark detection without changing the watermark generation process, we introduced an Entropy-based Text Watermark Detection (EWD) algorithm to accurately reflect the watermark level in the given text. Specifically, we argue that the influence of a token during detection should be proportional to its entropy. For example, the low-entropy tokens should have a limited impact on the detection result. Furthermore, in order to fully reflect this positive relation between token entropy and detection influence, we utilized a monotonically-increasing and continuous function to generate influence weights from token entropy. Then, inside the detector, the watermark level of a given text will be the weighted sum of all the green tokens. By creating a gap between the weights of high entropy tokens and low entropy tokens, we ensure that if the generation algorithm can hardly affect a low-entropy token, the state of that token will make minimal impact to the detected watermark result. The acquisition of token weight is an independent and training-free process given the whole text entropy, avoiding the trouble of a human dataset or any manual operations. Furthermore, the proposed detection scheme can be applied to machine-generated texts of all entropy levels. There is no need to switch between different detectors facing various kinds of input texts.
In the experiment, we evaluated the detection accuracy of EWD in low-entropy code generation task. We also test the versatility of our method in text of high entropy levels, such as document-to-text generation. Overall, our method outperformed baselinse methods including KGW and SWEET in detection accuracy. Besides, we included evaluation of our proposed EWD against watermark-removal attacks.
@@
  • We proposed an entropy-based text watermark detection method called EWD.
  • We provide a theoretical analysis of the detection accuracy in a low-entropy setting.
  • We showed that EWD not only improves the detection accuracy under the low-entropy scenario, but also provides similar performance as other methods in high-entropy texts.
@@ 2 Related Work
Watermark message can be embedded into existing text by performing a series of modifications, such as lexical substitutions Munyer and Zhong (2023); Topkara et al. (2006); Yang et al. (2022); Yoo et al. (2023a); Yang et al. (2023). Those methods typically replace words with their alternatives to place watermark. For example, Topkara et al. (2006) used WordNetFellbaum (1998) as his dictionary while Munyer and Zhong (2023) utilized a pretrained Word2Vec model to select the alternative words. To ensure that the repalced words would cause minimal changes to the text semantics, Yang et al. (2022) proposed a BERT-based infill model to generate replacement with regard to the sentence meaning. To improve the robustness of the algorithm against possible corruptions, Yoo et al. (2023a) fine-tuned the infill model to generate consistent word distribution given several corrupted texts of the same origin. However, the level of modification to an existing text is often limited to protect text quality, thus restricting the capacity and effectiveness of lexical-based watermarks.
@@
Watermark embedding can also be conducted during the LLM generation phase, either by modifying the token sampling or model logits. Christ et al. (2023) introduced a technique that incorporates watermark message by predefining a random number sequence for sampling the tokens. To enhance resistance to text modifications, Kuditipudi et al. (2023) utilized the Levenshtein distance in matching the text and numbers. For the latter, the most representative one that modifies the model logits is proposed by Kirchenbauer et al. (2023a) based on previous tokens during text generation. There are a number of works proposed on top of this algorithm with enhanced performance in payload Yoo et al. (2023b); Wang et al. (2023), robustness Zhao et al. (2023); Liu et al. (2023b); Kirchenbauer et al. (2023b); Ren et al. (2023) and unforgeabilityLiu et al. (2023a); Fairoze et al. (2023). For example, Yoo et al. (2023b) extends the method to a multi-bit setting.Wang et al. (2023) proposed to utilize a proxy LLM for multi-bit message embedding. To increase robustness against editing attacks, Zhao et al. (2023) proposed to set a universal green list during the generation process. Liu et al. (2023b) proposed a different approach which utilizes the semantic meaning of previous generated text to determine the green lists. Also, to enhance the unforgeability of text watermark, Liu et al. (2023a) introduced a neural network for text generation and detection. In this work, we focus only on the detection of low-entropy texts. Lee et al. (2023) designed a selective watermark generation and detection approach for low-entropy code generation. However, it requires a human code dataset to help manually determine the entropy threshold.
@@ 3 Preliminaries
3.1 LLM
Here we introduce the necessary concepts in this paper. A large language model (LLM) takes a prompt as input and outputs subsequent tokens as the corresponding response. Specifically, we denote the initial prompt as $x^{prompt}$. Then, at the $l$-th step, the input to the LLM is the $x^{prompt}$ and the already-generated tokens $T_{:(l-1)}$. The LLM would generate a distribution logits $P_{LLM}(x^{prompt},T_{:(l-1)})$, consisting of $P_{LLM}(v|x^{prompt},T_{:(l-1)})$ over each token $v$ of its model vocabulary based on the joint input.
@@ 3.2 Text Watermarking
Text watermarking is to embed a piece of message into the given text that can be decoded for the purpose of identification. Here we include a specific branch of it, defined as Watermark for LLMs by Liu et al. (2024), specifically the ones concerning Logits Generation. The general watermarking scheme used in this paper is called KGW Kirchenbauer et al. (2023a), which is to modify the logits generation process of an original LLM $M$ to acquire the watermarked LLM, denoted as $\tilde{M}$. The modification on the original logits by adding a small value to the logits of green-list tokens will result in LLM’s preference on those tokens in the output text. Thus such modification could likely produce a green token. The detection algorithm is to calculate the watermark level using the following equation
$z=(|s|_{G}-\gamma T)/\sqrt{\gamma(1-\gamma)T},$ (1)
where $|s|_{G}$ is the number of green tokens and $\gamma$ is the green-list ratio in the model vocabulary. The algorithm would output 1 if the value is significantly greater than a given threshold, and 0 otherwise. In this work, we only concern the watermark detection process.
@@ 3.3 Token Entropy and Low-entropy Scenario
For the logits during each generation step, its degree of spread-out can be measured by calculating the entropy $SE(P_{LLM}(x^{prompt},T_{:(l-1)}))$. In this paper, we follow the spike entropy from Kirchenbauer et al. (2023a),
$SE(k)=\sum_{v}{\frac{p_{v}}{1+zp_{v}}},$ (2)
where $p$ is the logits when sampling the token $k$ and $z$ is a scalar. The entropy value achieves minimal if the logits concentrates on a single token and maximal given a uniform distribution.The watermark generation process is more effective on a high-entropy token where the logits distribution is nearly uniform, so that the modification could produce a green-list token easily. On the contrary, a low-entropy token is almost deterministic and hard to be modified. Therefore, the green-list ratio for a low-entropy token is much lower, resulting in fewer green tokens in the output. A low-entropy scenario refers to the LLM-generated output consisting of a great number of low-entropy tokens. Here we denote the low-entropy token ratio to the whole text as $\varepsilon$.
4 Method
In this section, we will first explain the motivation by analysing the watermark effect on low-entropy tokens and traditional detection methods, then provide a detailed introduction to the proposed EWD.
@@ 4.1 Motivation
In KGW Kirchenbauer et al. (2023a)’s watermark generation method, a positive constant is added to the logits value of all green-list tokens so that the final output would prefer such tokens. The distribution level of the logits, represented by the entropy, matters significantly during this process. When the logits resemble a uniform distribution, meaning that the logits value among tokens are very close, this logits modification would very likely produce a green-list token. An average-high-entropy text would then consist of many green tokens and output a high z-score according to Eq.1. However, when the logits concentrate on one single token, meaning that it is almost deterministic to output that token, this watermark embedding process would hardly affect the token sampling and produce a green token with a much lower probability close to the green list ratio $\gamma$. As a result, the number of green tokens and the consequence detection z-score will be much lower for a high-$\varepsilon$ text, for example, the watermarked code in figure 1. This is because to the detector, tokens are of the same influence, no matter how difficult it is for the logits modification to affect them.
4.2 Proposed Method
In figure 1, we can observe that a high-entropy token is easier to be affected by the modified logits, and has a larger probability of being sampled from the green list. Since a low-entropy token is hard to affect, then whether it is sampled from the green list or not should have a limited influence on the overall watermark level. Specifically, we argued that when an easy-to-affect (high-entropy) token is not green, it should contribute more to showing that the text $T$ is not watermarked than a hard-to-affect (low-entropy) token. That is to say, The influence of a red token to yielding a negative detection result, $i.e.$, the given text is not watermarked, should be proportional to its entropy, and same for the influence of a green token to yielding a positive detection result:
$I(Detect(T)=\textbf{1}(t\in G)|t)\propto SE(t).$ (3)
When reflected in the real world, this influence index is represented by the token weight. Token with a larger spike entropy, meaning that it is easier to be affected, will be given a larger influence weight. On the contrary, a low-entropy token will be given a much lower weight close to 0. As the sum of all green token weights, $|s|_{G}$ will receive a smaller numeric loss when observing a red low-entropy token than a red high-entropy token, and a larger gain when observing a green high-entropy token than a green low-entropy token.

### v4

@@ Abstract
Text watermarking algorithms for large language models (LLMs) can effectively identify machine-generated texts by embedding and detecting hidden features in the text. Although the current text watermarking algorithms perform well in most high-entropy scenarios, its performance in low-entropy scenarios still needs to be improved. In this work, we opine that the influence of token entropy should be fully considered in the watermark detection process, $i.e.$, the weight of each token during watermark detection should be customized according to its entropy, rather than setting the weights of all tokens to the same value as in previous methods. Specifically, we propose Entropy-based Text Watermarking Detection (EWD) that gives higher-entropy tokens higher influence weights during watermark detection, so as to better reflect the degree of watermarking. Furthermore, the proposed detection process is training-free and fully automated. From the experiments, we demonstrate that our EWD can achieve better detection performance in low-entropy scenarios, and our method is also general and can be applied to texts with different entropy distributions. Our code and data is available [^1] https://github.com/luyijian3/EWD. Additionally, our algorithm could be accessed through MarkLLM Pan et al. (2024) [^2] https://github.com/THU-BPM/MarkLLM.
@@
Yijian LU1, Aiwei Liu2 , Dianzhi Yu1 , Jingjing Li1 , Irwin King1thanks: Corresponding Author. 1The Chinese University of Hong Kong 2Tsinghua University luyijian@link.cuhk.edu.hk, liuaw20@mails.tsinghua.edu.cn dianzhi.yu@link.cuhk.edu.hk, lijj@link.cuhk.edu.hk, king@cse.cuhk.edu.hk
@@ 1 Introduction
Figure 1: This figure shows that compared with watermarked texts with mostly high-entropy tokens, watermarked codes with mostly low-entropy tokens see significantly less green tokens, resulting in a small detection z-score. Furthermore, on the bottom of the figure, we demonstrate that the green token ratio in tokens decreases as their entropy decreases.
@@
The rapid advancements in large language models (LLMs) have enabled them to generate high-quality outputs indistinguishable from humans, and also achieve better performance on various real-world generation tasks. However, this gives rise to the potential risk of misuse. Specifically, LLM-generated assignments, including essays and codes, pose a threat to academic integrity. Misusing LLMs to generate fake news can also lead to negative societal consequences Megías et al. (2021). Therefore, algorithms capable of effectively identifying machine-generated contents have become essential.
Text watermarking algorithms can effectively alleviate the problem of LLM misuse by embedding and detecting hidden features in LLM-generated contents. For example, the watermarking algorithm proposed by Kirchenbauer et al. (2023a) (KGW) divides the LLM’s model vocabulary into two lists, green and red, and increases the logits value of tokens in the green list by a positive constant (watermark strength). The modified logits will bias the LLM to sample most tokens from the green list. As a result, the watermarked text containing a dominant number of green tokens will be recognized by the watermark detector while human texts will not. However, the successful watermark generation is largely tied to the token entropy during text generation. Under the watermark algorithm, the probability of sampling a token from the green list is proportional to its entropy, meaning that low-entropy tokens would produce far less green tokens than high-entropy tokens, as illustrated by Figure 1. In that case, green tokens will occupy only half of a watermarked low-entropy text, resulting in a smaller detection score and being mistakenly labeled as human-written. Furthermore, to ensure the text quality of output contents, the watermark strength in the generator cannot be excessively enlarged to forcefully modify the logits and sample more tokens from the green list Tu et al. (2023).
@@
Consequently, a crucial problem is how to improve the watermark detector so that low-entropy watermarked texts with a limited number of green tokens can be successfully identified. Lee et al. (2023) propose a selective watermark detector named SWEET to better reflect the watermark level in low-entropy texts. Specifically, they only detect tokens that exceed a certain entropy threshold, thus excluding tokens that are difficult to be affected by the watermark generator. However, this method itself has three weaknesses. First, their method requires a manually-set entropy threshold by analyzing samples of code written by humans. Second, their method also does not consider the distribution of token entropy, that is, their method treats all tokens that exceed the threshold as the same, while in fact the entropy of these tokens may vastly vary. Third, the performance gain using the SWEET detector is limited and there is still significant room for further advancement.
@@
To better address the low-entropy watermarking detection problem than previous methods, we introduce an Entropy-based Text Watermark Detection (EWD) algorithm to fully consider the role of entropy in the detection of a given text. Specifically, we propose that the influence of a token during detection should be proportional to its entropy. Furthermore, in order to fully reflect this positive relation between token entropy and detection influence, we utilize a monotonically-increasing and continuous function to generate influence weights from token entropy. By creating a gap between the weights of high entropy tokens and low entropy tokens, we ensure that if the watermark generation algorithm can hardly affect a low-entropy token, the state of that token will make minimal impact to the detection result. Overall, our method has three major benefits. First, the method saves the trouble of altering the watermark generation process. Second, compared to SWEET, our method is more efficient as the acquisition of token weights is an independent and training-free process, avoiding the trouble of a human dataset or any manual operations. Third, the proposed detection scheme can be applied to machine-generated texts of all entropy levels, not solely low-entropy texts.
From the experiments, we demonstrate that the detection accuracy of EWD in low-entropy scenarios surpasses existing baselines. Furthermore, EWD exhibits versatility in detecting texts of high entropy levels, as well as robustness against the back-translation watermark-removal attack.
@@
  • We propose an entropy-based text watermarking detection method called EWD.
  • We provide a theoretical analysis of the detection accuracy in the low-entropy setting.
  • We show that EWD not only improves the detection accuracy under the low-entropy scenario, but also provides similar performance as other methods in high-entropy texts.
@@ 2 Related Work
In this section, we aim to review existing works on text watermarking algorithms. Overall, text watermarking can be divided into two different categories: watermarking for existing texts and watermarking for LLMs Liu et al. (2023).
@@
Text Watermarking for Existing Texts. Watermark message can be embedded into existing text by performing a series of modifications, such as lexical substitutions Munyer and Zhong (2023); Topkara et al. (2006); Yang et al. (2022); Yoo et al. (2023a); Yang et al. (2023). Those methods typically replace words with their alternatives to place watermark. For example, Topkara et al. (2006) used WordNetFellbaum (1998) as his dictionary while Munyer and Zhong (2023) utilized a pretrained Word2Vec model to select the alternative words. To ensure that the repalced words would cause minimal changes to the text semantics, Yang et al. (2022) proposed a BERT-based infill model to generate replacement with regard to the sentence meaning. To improve the robustness of the algorithm against possible corruptions, Yoo et al. (2023a) fine-tuned the infill model to generate consistent word distribution given several corrupted texts of the same origin. However, the level of modification to an existing text is often limited to protect text quality in the watermarked output. Hence, watermarking for existing text is less favored due to its limited capacity and effectiveness.
@@
Text Watermarking for LLMs. Watermark embedding can also be conducted during the LLM generation phase, either by modifying the token sampling or model logits. Christ et al. (2023) introduced a technique that incorporates watermark message by predefining a random number sequence for sampling the tokens. To enhance resistance to text modifications, Kuditipudi et al. (2023) utilized the Levenshtein distance in matching the text and numbers. For the latter, the most representative one that modifies the model logits is proposed by Kirchenbauer et al. (2023a) based on previous tokens during text generation. There are a number of works proposed on top of this algorithm with enhanced performance in payload Yoo et al. (2023b); Wang et al. (2023), robustness Zhao et al. (2023); Liu et al. (2024b); Kirchenbauer et al. (2023b); Ren et al. (2023); He et al. (2024) and unforgeability Liu et al. (2024a); Fairoze et al. (2023). For example, Yoo et al. (2023b) extends the method to a multi-bit setting.Wang et al. (2023) proposed to utilize a proxy LLM for multi-bit message embedding. To increase robustness against editing attacks, Zhao et al. (2023) proposed to set a universal green list during the generation process. Liu et al. (2024b) proposed a different approach called SIR which utilizes the semantic meaning of previous generated text to determine the green lists. Inspired by Liu et al. (2024b), He et al. (2024) firstly proposed the Crosslingual Watermark Removal Attack (CWRA), then offered X-SIR as a defense method against CWRA. Also, to enhance the unforgeability of text watermark, Liu et al. (2024a) introduced a neural network for text generation and detection. In this work, we focus only on the detection of low-entropy texts. Lee et al. (2023) designed a selective watermark generation and detection approach for low-entropy code generation. However, it requires a human code dataset to help manually determine the entropy threshold. To sum up, there lacks an effective method to improve the low-entropy detection problem.
@@ 3 Preliminaries
3.1 Text Generation Process of LLMs
Here we introduce the necessary concepts used in this paper. A LLM, $M$ takes a prompt as input and outputs subsequent tokens as the corresponding response. Specifically, we denote the initial prompt as $x^{prompt}$. Then, at the $l$-th step, the input to the LLM is the $x^{prompt}$ and the already-generated tokens $T_{:(l-1)}$. The LLM would generate a distribution logits $P_{LLM}(x^{prompt},T_{:(l-1)})$, consisting of $P_{LLM}(v|x^{prompt},T_{:(l-1)})$ over each token $v$ of its model vocabulary based on the joint input.
@@ 3.2 Text Watermarking
The general watermarking scheme used in this paper is called KGW, which is to modify the logits generation process of an original LLM $M$ to acquire the watermarked LLM, denoted as $\tilde{M}$. The modification on the original logits by adding a small value (the watermark strength $\delta$) to the logits of green-list tokens will result in LLM’s preference on those tokens in the output text. Thus such modification could likely produce a green token.
The detection algorithm is to calculate the watermark level using the following equation,
$z=(|s|_{G}-\gamma|T|)/\sqrt{\gamma(1-\gamma)|T|},$ (1)
where $|s|_{G}$ is the number of green tokens in text $T$ and $\gamma$ is the green-list ratio in the model vocabulary. The algorithm would output $Detect(T)=1$ if the value is greater than a given threshold, and $Detect(T)=0$ otherwise.
@@ 3.3 Token Entropy and Low-entropy Scenario
The entropy of a token is measured by the degree of spread-out of the logits generated by the LLM during the sampling process of that token. In this paper, we follow the spike entropy from KGW,
$SE(k)=\sum_{v}{\frac{p_{v}}{1+\tau p_{v}}},$ (2)
where $p_{v}$ is the logits value of each token $v$ in the vocabulary when sampling the token $k$, and $\tau$ is a scalar. The entropy value reaches the minimal value if the logits concentrates on a single token and the maximal value given a uniform distribution.
The watermark generation process is more effective on a high-entropy token where the logits distribution is nearly uniform, so that the modification could produce a green-list token easily. On the other hand, a low-entropy token is almost deterministic and hard to be affected. Therefore, the green-list ratio for a low-entropy token is much lower, resulting in fewer green tokens in the output. Therefore, low-entropy watermarked texts pose threats on the detection accuracy as they are easily identified as human-written.
4 Proposed Method
In this section, we will first explain the motivation by analysing the watermark generation effect on low-entropy tokens, then provide a detailed explanation of the proposed EWD.
@@ 4.1 Motivation
Entropy plays an important role in the watermark generation process, yet its role in the detection process has been underestimated. In the KGW generator, a positive constant is added to the logits value of all green-list tokens. The distribution level of the logits, represented by the entropy, matters significantly during this process. When the logits resembles a uniform distribution, this logits modification would very likely produce a green-list token. However, when the logits concentrates on one single token, this watermark embedding process would hardly affect the token sampling and produce a green token with a much lower probability close to the green list ratio $\gamma$. Since entropy determines the watermark generation process, then it should also affect the watermark detection. However, in the KGW detector, all tokens are of the same influence, no matter how their entropy varies.
4.2 Entropy-based Text Watermarking Detection
Our proposed EWD aims to fully consider the role of entropy in the watermark detection process. Specifically, we firstly propose a positive influence-entropy relation, then utilize entropy for detection weight customization.
The influence of a token during detection should be proportional to its entropy. From previous sections, we explain why a high-entropy token is easier to be affected by the watermark generator, as well as the situation for a low-entropy token. Since a low-entropy token is hard to be affected, then whether it is sampled from the green list or not should have a limited influence on the overall watermark level. Similarly, when an easy-to-affect (high-entropy) token is not green, it should contribute more to showing that the text $T$ is not watermarked. Therefore, we introduce a positive influence-entropy relation to determine how much a token $t$ can influence the detection result:
$I(Detect(T)=\textbf{1}(t\in G)|t)\propto SE(t),$ (3)
where $T$ is the given text and $G$ is the green list.
The detection weight of a token should be customized based on its entropy to fully reflect the positive influence-entropy relation. The influence index in Eq. (3) is represented by the detection weight of each token. However, the detection weight assigning of each token in SWEET fails to reflect the positive influence-entropy relation. This is because that inside the two groups of above-threshold tokens and below-threshold tokens, the weights of tokens with different entropy are identical. Therefore, to fully reflect the influence-entropy relation, we utilize a monotonically-increasing and continuous function $f$ to customize weights from token entropy. The weight $W(t)$ of a token $t$ can be illustrated as follows:
$W(t)=f(SE(t)-C_{0}),$ (4)
where $C_{0}$ is the minimal value of the spike entropy to normalize the entropy input before computing the weight. Using the entropy-based weight customization, even a sight increase in the token entropy will result in a increased detection weight. Table 1 summarizes the weight features of different methods and their reflection of the proposed influence-entropy relation.
Table 1: Features of detection weights by different methods and how they can reflect the proposed positive influence-entropy relation.
Methods Detection Weight
Reflection of
Eq. (3)
KGW Constant None
SWEET Binary Partial
EWD Customized Full